A stock has an expected return of 10 percent, its beta is .9 and the risk-free...

  1. A stock has an expected return of 10 percent, its beta is .9 and the risk-free rate is 6 percent. What must the expected return on the market be?

PLEASE GIVE ME A FULL EXPLANATION WITH THE FORMULAS. THANK YOU
Answer #1

expected return=risk free rate+beta*(market rate-risk free rate)

10=6+0.9*(market rate-6)

(10-6)=0.9*(market rate-6)

market rate=[(10-6)/0.9]+6

=10.44%(Approx).